### v1

@@ 3.2 Geometry-guided camera-to-BEV transformation
With the extracted image feature $\boldsymbol{F}_{img}\in\mathbb{R}^{U\times V\times C}$, we infer a BEV feature map $\boldsymbol{F}_{bev}\in\mathbb{R}^{L\times D\times C}$ in a $L\times D$ grid map on the x-y plane of the global coordinate system. Following the BEV perception methods (Philion and Fidler, 2020; Liu et al., 2023), for each pixel $(u,v)$ on the image, we sample D discrete points $\{(u,d_{i},v)\in\mathbb{R}^{3}|d_{i}=d_{0}+i\Delta,i\in\{1,2,...,D\}\}$ among the camera ray and predict the associated depth distribution $\boldsymbol{\alpha}_{u,v}\in\mathbb{R}^{D}$. Then we scatter the image feature to the points with the corresponding probability, which generates the feature point cloud $\boldsymbol{F}_{pc}\in\mathbb{R}^{\text{UDV}\times C}$ of size UDV. For each point $(u,d_{i},v)$, the associated feature $\boldsymbol{F}_{pc}(u,d_{i},v)\in\mathbb{R}^{C}$ is defined as the rescaled feature vector of pixel $(u,v)$:
@@
Since the image is gravity-aligned, each column corresponds to a triangular "slice" in the frustum (Lentsch et al., 2023), and the associated feature points should lie within the same rectangular region extending from the camera in the BEV polar coordinate system. Therefore, we assemble the feature point cloud of each column into a polar strip on the plane, resulting in a $D\times V$ polar feature map $\boldsymbol{F}_{bev}$. For each polar grid $(d,v)$, the feature projection could be formulated as:
$\boldsymbol{F}_{bev}(d,v)=\sum_{u=1}^{U}\alpha_{u,d,v}\boldsymbol{F}_{pc}(u,d,v),$ (2)
@@
Depth distribution adapter: Although simple, unsupervised monocular depth estimation presents significant challenges. While the depth value might be available from other sensors (e.g., LiDAR or RGB-D camera) for supervision, it is a non-trivial assumption and hard to generalize to novel devices and unseen environments. Furthermore, for each pixel, converting the depth value $d\in(0,+\infty)$ into the discrete depth distribution $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ is also an ill-posed problem.
Fortunately, Depth Anything (Yang et al., 2024) offers a zero-shot relative depth estimation model, where the output normalized disparity $\overline{t}\in[0,1]$ provides powerful prior information about the relative distance relationship for view transformation. The normalized disparity $\overline{t}$ is a scale- and shift-invariant, numerically stable version of the depth $d$. During supervised training, the authors first transform the depth value $d$ into the disparity $t$ by $t=1/d$ and then normalize it to the $\overline{t}$ for each image:
$\overline{t}=\frac{t-\text{min}(t)}{\text{max}(t)-\text{min}(t)},$ (3)
where $\text{max}(t)$ and $\text{min}(t)$ are the maximal and minimal disparity values within the image. However, both $\text{max}(t)$ and $\text{min}(t)$ are unavailable during inference, and the normalized disparity $\overline{t}$ cannot be reversed to the depth $d$ or the depth distribution $\boldsymbol{\alpha}$ for the view transformation.
Considering the above issue, we design the depth distribution adapter (DDA) to distill the prior knowledge from the Depth Anything (Yang et al., 2024) model into the camera-to-BEV transformation, as shown in Fig. 4. With the predicted depth distribution $\boldsymbol{\alpha}_{u,v}\in\mathbb{R}^{D}$ for each pixel $(u,v)$, the depth value $d_{u,v}$ is formatted as the weighted average value of the depth distribution:
Then we transform the depth estimation $d_{u,v}$ to normalized disparity $\overline{t}_{u,v}$ according to Eq. (3). With the pseudo ground truth (GT) disparity value $\hat{t}_{u,v}$ from the decoder of the Depth Anything model, the depth distribution loss $\mathcal{L}_{depth}$ is defined as the mean absolute error:
$\mathcal{L}_{depth}=\frac{1}{UV}\sum_{u=1}^{U}\sum_{v=1}^{V}|\overline{t}_{u,v}-\hat{t}_{u,v}|.$ (5)
By aligning the disparity predictions with Depth Anything, the DDA module effectively incorporates geometric prior from the foundation model into the framework, enabling it to understand the topology of surrounding objects. During inference, we discard the disparity supervision and directly utilize the depth distribution $\boldsymbol{\alpha}$ for the camera-to-BEV transform.
Figure 4: Illustration of the depth distribution adapter. For each pixel, we predict the depth distribution $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ and scatter the image feature $\boldsymbol{F}_{img}$ to the BEV feature $\boldsymbol{F}_{bev}$ by the corresponding depth distribution $\boldsymbol{\alpha}$. During training, the depth distribution prediction is transferred to the depth map and supervised with the Depth Anything.
@@ 3.3 Semantic-guided neural localization
Since OSM data provides semantic and geographic information about stationary objects, the key to successful localization lies in constructing the BEV feature $\boldsymbol{F}_{bev}$ with consistent semantics and topology to those of the OSM data. Explicit feature registration at the semantic level also aligns with how humans localize themselves (Epstein and Vass, 2014). However, previous I2O visual localization approaches largely overlooked the semantic consistency of images and OSM data. OrienterNet (Sarlin et al., 2023) exploited the network for pose estimation in an end-to-end way. However, pose supervision only makes it challenging for the model to fully understand the surroundings. Concurrent approach MaplocNet (Wu et al., 2024) pointed out that the semantic labels from the HD maps could improve the comprehension ability of the image model and substantially enhance visual localization accuracy. However, these semantic labels are not freely accessible and limit the method’s generalizability. To address the challenge, we introduce an auxiliary semantic alignment task to improve the scene understanding capability of the image model. The semantic alignment task fully utilizes the semantic embeddings $\mathcal{M}_{sem}$ as guidance and is generally free. During the training phase, for each grid $(l,d_{i})$ on the BEV feature $\boldsymbol{F}_{bev}\in\mathbb{R}^{L\times D\times C}$, we find the corresponding area on the map tile $\mathcal{M}$ and the associated semantic embedding vector with the GT pose $\hat{\boldsymbol{p}}$. Then, we directly align the BEV features $\boldsymbol{F}_{bev}$ with the corresponding semantic embeddings of the OSM data. The semantic alignment loss $\mathcal{L}_{sem}$ is defined as:
$\mathcal{L}_{sem}=\frac{1}{LD}\sum_{l=1}^{L}\sum_{i=1}^{D}\|C_{1\times 1}(\boldsymbol{F}_{bev}[l,d_{i}])-\boldsymbol{M}_{sem}[\hat{\Gamma}(l,d_{i})]\|_{2}.$ (6)
$\hat{\Gamma}$ is the transformation matrix of GT pose $\hat{\boldsymbol{p}}$ and $C_{1\times 1}$ is a $1\times 1$ convolution layer to adjust the channel dimension of the BEV features.
@@
With the semantic-rich BEV features $F_{bev}$ and map features $F_{map}$, the neural localization is responsible for estimating the 3-DoF pose $\boldsymbol{p}=(x,y,\theta)$ of the image. The most straightforward strategy is, fusing the BEV features and map features and estimating the pose with a small regressor. Unfortunately, in our pilot studies, the network failed to converge with the regression design after trial and error, which contradicts the observation in MaplocNet (Wu et al., 2024). We speculate that this is due to the limited field of view (FoV) of a single image, which results in the BEV feature $\boldsymbol{F}_{bev}$ containing partial information. Therefore, directly learning the pose vector is challenging. To simplify the problem, we discretely sample the continuous 3-D pose space at each map grid and $K$ rotations with regular intervals, resulting in a $H\times W\times K$ volume of pose candidates in the 3-DoF pose space. For each pose candidate $\boldsymbol{p}_{h,w,k}=(x_{h},y_{w},\theta_{k})$, we project the BEV feature $\boldsymbol{F}_{bev}$ onto the map plane and calculate the matching score in the feature space :
$\boldsymbol{S}_{h,w,k}=\frac{1}{LD}\sum_{l=1}^{L}\sum_{i=1}^{D}\boldsymbol{F}_{map}(\Gamma_{h,w,k}(l,d_{i}))\odot\boldsymbol{F}_{bev}(l,d_{i}),$ (7)
@@
where $\Gamma_{h,w,k}$ is the transformation matrix for pose candidate $\boldsymbol{p}_{h,w,k}$, and $\odot$ denotes the inner product operation. Eq. 7 benefits from an efficient implementation by performing a single convolution as a batched multiplication in the Fourier domain (Barsan et al., 2018). Finally, the scoring volume $\boldsymbol{S}$ is normalized to the probability volume $\boldsymbol{P}$ with the softmax operation.
@@
The pose classification loss $\mathcal{L}_{pose}$ is defined as a binary cross-entropy loss to maximize the probability of the GT pose $\hat{\boldsymbol{p}}$:
@@
where the $\hat{P}$ is the probability volume of the GT pose, which is concentrated at a single point $\boldsymbol{\hat{p}}=(\hat{x},\hat{y},\hat{\theta})$ in the pose space.
Overall, the joint loss $\mathcal{L}$ is defined as a weighted combination of pose classification loss $\mathcal{L}_{pose}$, depth distribution loss $\mathcal{L}_{depth}$ and semantic alignment loss $\mathcal{L}_{sem}$:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{pose}+\lambda_{2}\mathcal{L}_{depth}+\lambda_{3}\mathcal{L}_{sem},$ (9)
@@
During the inference, we estimate the pose $\boldsymbol{p}^{*}$ as the one corresponding to the peak in the probability volume $\boldsymbol{P}$:
@@ 4 Cross-area and cross-condition validation benchmark
OrienterNet (Sarlin et al., 2023) has constructed the Mapillary geo-localization (MGL) dataset for the I2O visual localization task, containing 826 K images for training and 2 K images for validation. The training-evaluation split follows the "same-area" protocol, where each location’s images are disjointed into training and validation sets. Besides, the distribution of cameras, motions, viewing conditions, and visual features of the validation set is identical to the training set. Such same-area and same-condition setup may work well for localization within a small, local area over a short time span, it is not appropriate for assessing the performance of I2O localization methods across different countries with VGI data collected at various conditions.
To address this concern, we propose a novel validation benchmark for cross-area and cross-condition validation, called the CC validation benchmark in the following. Table. 1 lists details of the CC validation benchmark, which contains 7164 images from 4 cities on the Mapillary platform (Mapillary, 2024). These images were collected with diverse devices, by different users, with varying motion patterns and under various external conditions, as shown in Fig. 5 and Fig. 6. They have never been seen before and are located hundreds of miles away from the nearest area in the MGL dataset. Each image is captured by cameras handheld or mounted on helmets, UGVs, or vehicles. For each 360 panorama, we split it into $4\times 90^{\circ}$ FoV perspective images with a random offset and resample them to a size of $512\times 512$ pixels. We query the OSM data for the corresponding area from the official website (OSM, 2024) and the data pre-processing is shown in Fig. 2. The validation scenes in the CC validation benchmark cover a broad range of urban scenes, including downtown, suburban, overpass regions, and pedestrian streets. We believe the diverse set of scenes, captured by various cameras, users, platforms, and under different external conditions, makes the CC benchmark more challenging and suitable for comprehensive evaluation.
@@
Munich 2156 Trimble mx50 Handheld
@@ 5.1 Experimental setup
We evaluate the performance of our OSMLoc on three datasets: the MGL dataset, the CC validation benchmark, and the KITTI (Geiger et al., 2012) dataset. Our model is trained on the MGL dataset and directly validated on the CC validation benchmark and KITTI dataset without fine-tuning.
@@
  • CC validation benchmark: consists of 14 long sequences with 7164 images from Detroit, Munich, Taipei, and Brisbane captured between 2017 to 2024. The regions have never been seen in the MGL dataset. More details can be found in Sec. 4.
  • KITTI dataset: provides the stereo images captured by a moving vehicle along different trajectories, primarily used for autonomous driving. We use the KITTI dataset to evaluate the generalization capability of the proposed method across various image-capturing platforms. The GT poses were collected from the integrated navigation system. Shi and Li (2022) split the KITTI dataset into the Train, Test1 and Test2 sets for I2A visual localization task. The Test2 set was collected in a different area, with no overlap of the MGL dataset or the KITTI Train set. Following the OrienterNet (Sarlin et al., 2023), we select the Test2 as the validation set, assuming that an accurate initial pose sampled within ±$\pm 20\text{m}$ and ±$\pm 10^{\circ}$ is available. Only the pose candidates that fall within the range are considered during inference.
Baseline methods: Due to the limited focus on the single-image I2O visual localization, OrieneterNet (Sarlin et al., 2023) is currently the only open-sourced comparable method in the field. We introduce variants of I2A visual localization methods (Shi and Li, 2022; Xia et al., 2022) as baselines for further comparison.
  • OrienterNet (Sarlin et al., 2023): is the pioneering end-to-end I2O visual localization method for single image. We use the officially released code for implementation and retrain the model with the same settings, except for setting the batch size to 4.
  • Retrieval (Xia et al., 2022): replaces the BEV inference and matching with a correlation between the neural map and a global image embedding. As the original implementation predicts a 2-DoF position and ignores the orientation, we predict the orientation by considering four neural maps for the North-South-East-West directions. This formulation also regresses a probability volume, similar to OrienterNet.
  • Refinement (Shi and Li, 2022): updates an initial pose by warping a satellite view to the image assuming that the scene is planar, at a fixed height, and gravity-aligned. We replace the satellite view with a rasterized OSM tile. This formulation requires an initial orientation (both during training and testing), which we sample within $45^{\circ}$ of the ground truth orientation angle.
@@
$\text{PR}=\frac{1}{M}\sum_{i=1}^{M}{\mathds{1}(\|(x_{i},y_{i})-(\hat{x}_{i},\hat{y}_{i})\|_{2}<\sigma_{p}),}$ (11)
$\text{OR}=\frac{1}{M}\sum_{i=1}^{M}\mathds{1}(|\theta-\hat{\theta}_{i}|<\sigma_{o}),$
where $M$ is the number of test frames and the threshold pairs $\boldsymbol{\sigma}=(\sigma_{p},\sigma_{o})$ are the maximal tolerance of the localization error. The APE and AOE are the mean absolute error between the predicted poses and GT poses:
$\text{APE}=\frac{1}{M}\sum_{i=1}^{M}\|(x_{i},y_{i})-(\hat{x}_{i},\hat{y}_{i})\|_{2},$ (12)
$\text{AOE}=\frac{1}{M}\sum_{i=1}^{M}|\theta_{i}-\hat{\theta}_{i}|.$
For the KITTI dataset, since position recall is closely related to the vehicle’s motion direction in autonomous driving scenes (Shi and Li, 2022), we separate the position recalls into two components: the perpendicular direction recall (latitudinal, LatR) and the parallel direction recall (longitudinal, LonR) to the viewing axis.

### v3

@@ 3.2 Geometry-guided camera-to-BEV transformation
With the extracted image feature $\boldsymbol{F}_{img}\in\mathbb{R}^{U\times V\times C}$, we infer a BEV feature map $\boldsymbol{F}_{bev}\in\mathbb{R}^{D\times L\times C}$ in a $D\times L$ grid map on the x-y plane of the global coordinate system. Following the BEV perception methods [56, 52], for each pixel $(u,v)$ on the image, we sample D discrete points $\{(u,d_{i},v)\in\mathbb{R}^{3}|d_{i}=d_{0}+i\Delta,i\in\{1,2,...,D\}\}$ among the camera ray and predict the associated depth distribution $\boldsymbol{\alpha}_{u,v}\in\mathbb{R}^{D}$. Then we scatter the image feature to the points with the corresponding probability, which generates the feature point cloud $\boldsymbol{F}_{pc}\in\mathbb{R}^{\text{UDV}\times C}$ of size UDV. For each point $(u,d_{i},v)$, the associated feature $\boldsymbol{F}_{pc}(u,d_{i},v)\in\mathbb{R}^{C}$ is defined as the rescaled feature vector of pixel $(u,v)$:
@@
With the predicted depth distribution $\boldsymbol{\alpha}_{u,v}\in\mathbb{R}^{D}$ for each pixel $(u,v)$, the depth value $d_{u,v}$ is formatted as the weighted average value of the depth distribution:
Since the image is gravity-aligned, each column corresponds to a triangular "slice" in the frustum [57], and the associated feature points should lie within the same rectangular region extending from the camera in the BEV polar coordinate system. Therefore, we assemble the feature point cloud of each column into a polar strip on the plane, resulting in a $D\times V$ polar feature map $\boldsymbol{F}_{bev}$. For each polar grid $(d_{i},v)$, the feature projection could be formulated as:
$\boldsymbol{F}_{bev}(d_{i},v)=\sum_{u=1}^{U}\alpha_{u,d_{i},v}\boldsymbol{F}_{pc}(u,d_{i},v),$ (3)
@@
Depth distribution adapter: Although conceptually straightforward, accurate depth distribution is generally unavailable for single-view images in real-world scenarios. External sensors such as LiDAR or RGB-D cameras may provide such information, but relying on them imposes a strong assumption that limits generalization to novel devices and unseen environments. Absolute depth estimators [58, 59] aim to recover depth from a single image; however, even state-of-the-art methods still exhibit substantial errors, often ranging from several to tens of meters. In contrast, relative depth estimation in a zero-shot manner provides a viable alternative by predicting normalized disparity $\overline{t}$, which preserves ordinal depth relationships between scene elements (e.g., "object A is closer than object B") and remains robust without requiring metric supervision or dataset-specific fine-tuning. We first describe the transformation relationships among depth, disparity, and normalized disparity here. Let the depth at pixel $(u,v)$ be denoted by $d_{u,v}$. The disparity is defined as its inverse, $t_{u,v}=1/d_{u,v}$, and the normalized disparity $\overline{t}_{u,v}$ for each image is given by:
$\overline{t}_{u,v}=\frac{t_{u,v}-\min(t)}{\max(t)-\min(t)},\quad\overline{t}_{u,v}\in(0,1),$ (4)
where $\max(t)$ and $\min(t)$ denote the extrema of disparity in the image. Intuitively, $\overline{t}_{u,v}$ can be viewed as a probability-like quantity that distributes pixels along a $(0,1)$ interval, thereby capturing the structural layout of the scene without requiring metric consistency across images.
With the pseudo ground-truth normalized disparity $\hat{t}_{u,v}$ generated from DAM decoder [26], we thus transform the predicted depth $d_{u,v}$ into normalized disparity $\overline{t}_{u,v}$ with Eq. (4) first, and then supervise the predicted $\overline{t}_{u,v}$ against $\hat{t}_{u,v}$ using an L1 loss on the image:
$\mathcal{L}_{dis}=\frac{1}{UV}\sum_{u=1}^{U}\sum_{v=1}^{V}\big{|}\overline{t}_{u,v}-\hat{t}_{u,v}\big{|},$ (5)
which encourages the network to learn the ordinal depth distribution while transferring geometric priors from DAM into our framework. For intuitive understanding, the entire process of DDA is depicted in Fig. 4.
Figure 4: Illustration of the depth distribution adapter. For each pixel, we predict the depth distribution $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ and scatter the image feature $\boldsymbol{F}_{img}$ to the BEV feature $\boldsymbol{F}_{bev}$ by the corresponding depth distribution $\boldsymbol{\alpha}$. During training, the depth distribution prediction is transferred to the normalized disparity map and supervised with the Depth Anything.
@@ 3.3 Semantic-guided neural localization
Semantic Alignment: Since OSM data provides semantic and geographic information about stationary objects, the key to successful localization lies in constructing the BEV feature $\boldsymbol{F}_{bev}$ with consistent semantics and topology to those of the OSM data. Explicit feature registration at the semantic level also aligns with how humans localize themselves [2]. However, previous I2O visual localization approaches largely overlooked the semantic consistency of images and OSM data. OrienterNet [21] exploited the network for pose estimation in an end-to-end way, whereas pose supervision only makes it challenging for the model to fully understand the surroundings. Concurrent approach MaplocNet [22] pointed out that the semantic labels from the HD maps could improve the comprehension ability of the image model and substantially enhance visual localization accuracy. However, these semantic labels are not freely accessible and limit the method’s generalizability.
[…]
$\mathcal{L}_{sem}=\frac{1}{DL}\sum_{i=1}^{D}\sum_{l=1}^{L}\big{\|}C_{1\times 1}(\boldsymbol{F}_{bev}[d_{i},l])-\boldsymbol{M}_{sem}[\hat{\Gamma}(d_{i},l)]\big{\|}_{2},$ (6)
where $C_{1\times 1}$ denotes a $1\times 1$ convolution layer for channel adjustment. The proposed semantic alignment task aligns BEV features with OSM semantic embeddings, enabling the image network to better perceive its surroundings in the BEV space. This enhanced perception benefits neural matching-based localization and comes essentially free of additional annotation.
@@
With the semantic-rich BEV features $F_{bev}$ and map features $F_{map}$, the neural localization is responsible for estimating the 3-DoF pose $\boldsymbol{p}=(x,y,\theta)$ of the image. The most straightforward strategy is, fusing the BEV features and map features and estimating the pose with a small regressor. Unfortunately, in our pilot studies, the network failed to converge with the regression design after trial and error, which contradicts the observation in MaplocNet [22]. We speculate that this is due to the limited field of view (FoV) of a single image, which results in the BEV feature $\boldsymbol{F}_{bev}$ containing partial information. Therefore, directly learning the pose vector is challenging. To simplify the problem, we discretely sample the continuous 3-D pose space at each map grid and $K$ rotations with regular intervals, resulting in a $H\times W\times K$ volume of pose candidates in the 3-DoF pose space. For each pose candidate $\boldsymbol{p}_{h,w,k}=(x_{h},y_{w},\theta_{k})$, we project the BEV feature $\boldsymbol{F}_{bev}$ onto the map plane and calculate the matching score in the feature space :
$\boldsymbol{S}_{h,w,k}=\frac{1}{DL}\sum_{i=1}^{D}\sum_{l=1}^{L}\boldsymbol{F}_{map}(\Gamma_{h,w,k}(d_{i},l))\odot\boldsymbol{F}_{bev}(d_{i},l),$ (6)
@@
where $\Gamma_{h,w,k}$ is the transformation matrix for pose candidate $\boldsymbol{p}_{h,w,k}$, and $\odot$ denotes the inner product operation. Eq. (6) benefits from an efficient implementation by performing a single convolution as a batched multiplication in the Fourier domain [60]. Finally, the scoring volume $\boldsymbol{S}$ is normalized to the probability volume $\boldsymbol{P}$ with the softmax operation.
@@
The pose loss $\mathcal{L}_{pose}$ is defined as a binary cross-entropy loss to maximize the probability of the GT pose $\hat{\boldsymbol{p}}$:
@@
[…]
Overall, the joint loss $\mathcal{L}$ is defined as a weighted combination of pose loss $\mathcal{L}_{pose}$, disparity loss $\mathcal{L}_{dis}$ and semantic alignment loss $\mathcal{L}_{sem}$:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{pose}+\lambda_{2}\mathcal{L}_{dis}+\lambda_{3}\mathcal{L}_{sem},$ (6)
@@
During inference, we estimate the pose $\boldsymbol{p}^{*}$ as the one corresponding to the maximal value in the probability volume $\boldsymbol{P}$:
@@ 4 Cross-area and cross-condition validation benchmark
OrienterNet [21] has constructed the MGL dataset for the I2O visual localization task, containing 826 K images for training and 2 K images for validation. The training-evaluation split follows the "same-area" protocol, where each location’s images are disjointed into training and validation sets. Besides, the distribution of cameras, motions, viewing conditions, and visual features of the validation set is identical to the training set. Such same-area and same-condition setup may work well for localization within a small, local area over a short time span, it is not appropriate for assessing the performance of I2O localization methods across different countries with VGI data collected at various conditions.
To address this concern, we propose a novel validation benchmark for cross-area and cross-condition validation, called the CC validation benchmark in the following. Table. 1 lists details of the CC validation benchmark, which contains 7164 images from 4 cities on the Mapillary platform [61]. These images were collected with diverse devices, by different users, with varying motion patterns and under various external conditions, as shown in Fig. 5 and Fig. 6. They have never been seen before and are located hundreds of miles away from the nearest area in the MGL dataset. Each image is captured by cameras handheld or mounted on helmets, UGVs, or vehicles. For each 360 panorama, we split it into $4\times 90^{\circ}$ FoV perspective images with a random offset and resample them to the size of $512\times 512$ pixels. We query the OSM data for the corresponding area from the official website [62] and the data pre-processing is shown in Fig. 2. The validation scenes in the CC validation benchmark cover a broad range of urban scenes, including downtown, suburban, overpass regions, and pedestrian streets. We believe the diverse set of scenes, captured by various cameras, users, platforms, and under different external conditions, makes the CC benchmark more challenging and suitable for comprehensive evaluation.
@@
Munich 2156 Trimble mx50 Helmet
@@ 5.1 Experimental setup
We evaluate the performance of our OSMLoc on three datasets: the MGL dataset, the CC validation benchmark, and the KITTI [63] dataset. Our model is trained on the MGL dataset and directly validated on the CC validation benchmark and KITTI dataset without fine-tuning.
@@
  • CC benchmark: We construct the cross-area and cross-condition (CC) benchmark with 7,164 images from four unseen cities: Detroit, Munich, Taipei, and Brisbane, which are located hundreds of kilometers from the MGL training scenes. This benchmark is designed to assess the global generalization capability of I2O approaches.
  • KITTI dataset: KITTI [63] provides stereo images captured by a moving vehicle in Karlsruhe, Germany, which is utilized to demonstrate the application value of our OSMLoc in the autonomous driving field and various image-capturing platforms. Shi and Li [64] split the KITTI dataset into the Train, Test1 and Test2 sets for I2A visual localization task. The Test2 set was collected in a different area, with no overlap of the MGL dataset or the KITTI Train set. Following the OrienterNet [21], we select the Test2 as the validation set, assuming that an accurate initial pose sampled within $\pm 20\text{m}$ and $\pm 10^{\circ}$ is available. Only the pose candidates that fall within the range are considered during inference.
Baseline methods: Due to the limited focus on the single-image I2O visual localization, OrieneterNet [21] is currently the only open-sourced comparable method in the field. We adopt variants of I2A visual localization methods [64, 65, 32], along with the foundational model based method AnyLoc [54], as baselines for comparison.
  • OrienterNet [21]: is the pioneering end-to-end I2O visual localization method for single image. The image encoder is a ResNet101-FPN [66] pre-trained on ImageNet, while the map encoder is a VGG19-FPN [67] architecture trained from scratch. We use the officially released code for implementation and retrain the model with the same settings, except for setting the batch size to 4.
  • Retrieval [65]: replaces BEV inference and matching with a correlation between the neural map and a global image embedding. While the original implementation only estimates a 2-DoF position without orientation, we extend it by predicting orientation through four neural maps corresponding to the North, South, East, and West directions.
  • Refinement [64]: updates an initial pose by warping a satellite view to the image, assuming that the scene is planar, at a fixed height, and gravity-aligned. We replace the satellite image with a rasterized OSM tile. This method requires an initial orientation during both training and evaluation, which we sample within $45^{\circ}$ of the ground truth orientation angle.
  • SNAP [32]: is a ground image-to-aerial image localization method. Following the official implementation, we utilize a ResNet50-FPN network as the image encoder and a VGG19-FPN network as the map encoder. InfoNCE [68] loss is introduced to maximize the correct localization probability and minimize the mismatch probability.
  • AnyLoc [54]: is originally developed for visual place recognition and leverages the foundational model DINOv2 [25] to avoid re-training or fine-tuning. However, since OSM tiles contain only semantic information rather than color or texture cues, we observed that the foundation model fails to extract sufficiently discriminative features. To address this, we employ the DINOv2-S model as the image encoder and NetVLAD as the feature aggregator for end-to-end training, while adopting a VGG19-FPN architecture as the map encoder. To handle orientation, we further predict rotation by generating four neural maps corresponding to the North-South-East-West directions.
@@
$\text{PR}=\frac{1}{N}\sum_{i=1}^{N}{\mathds{1}(\|(x_{i},y_{i})-(\hat{x}_{i},\hat{y}_{i})\|_{2}<\sigma_{p}),}$ (6)
$\text{OR}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(|\theta-\hat{\theta}_{i}|<\sigma_{o}),$
where $N$ is the number of test frames and the threshold pairs $\boldsymbol{\sigma}=(\sigma_{p},\sigma_{o})$ are the maximal tolerance of the localization error. The APE and AOE are the mean absolute error between the predicted poses and GT poses:
$\text{APE}=\frac{1}{N}\sum_{i=1}^{N}\|(x_{i},y_{i})-(\hat{x}_{i},\hat{y}_{i})\|_{2},$ (6)
$\text{AOE}=\frac{1}{N}\sum_{i=1}^{N}|\theta_{i}-\hat{\theta}_{i}|.$
For the KITTI dataset, since position recall is closely related to the vehicle’s motion direction in autonomous driving scenes [64], we decompose the position recall and error into two components relative to the viewing axis: recall and absolute error in the perpendicular (latitudinal) direction (LatR and ALatE) and in the parallel (longitudinal) direction (LonR and ALonE).